OpenStudy (anonymous):

Determine how many, what type, and find the roots for f(x) = x4 + 21x2 − 100.

OpenStudy (hayleymeyer):

lets take t=x^2

OpenStudy (welshfella):

its degree 4 so how many roots will there be?

OpenStudy (anonymous):

4

OpenStudy (hayleymeyer):

so we have the equation t^2 +21t-100 =0

OpenStudy (welshfella):

yea

OpenStudy (hayleymeyer):

Looking at the discriminant we have discriminant = \[\frac{ -21\pm \sqrt{21^2-4(1)(-100)} }{ 2}\]

OpenStudy (welshfella):

- which can be factored

OpenStudy (welshfella):

no need this can be factored

OpenStudy (hayleymeyer):

simplifying the discriminant we get - t=4 or t= -25

OpenStudy (hayleymeyer):

we know that x^2 =t so \[x=\sqrt{t}\]

OpenStudy (anonymous):

ok

OpenStudy (hayleymeyer):

we have 2 values of t :) one is positive and one is negative :) what do u think can we put negative value of t in this function to get x- \[x=\sqrt{t}\]?

OpenStudy (anonymous):

idk to be honest

OpenStudy (welshfella):

t = 4 gives 2 values for x 2 and - 2

OpenStudy (hayleymeyer):

ok well can the square of any number be negative ?

OpenStudy (welshfella):

now the square roots of -25 are imaginary do you know what they are?

OpenStudy (welshfella):

yes - you introduce the operator i which stands for the square root of -1.

OpenStudy (anonymous):

5i?

OpenStudy (welshfella):

so sqrt -15 = -5i and 5 i

OpenStudy (welshfella):

*sqrt -25

OpenStudy (welshfella):

and theres your 4 roots 2, -2 , 5i and -5i

OpenStudy (hayleymeyer):

well \[\sqrt{-25} \] does not exists cause -25 is negative so we r left with t=4 puttin t=4 in the equation \[x=\sqrt{t}\]we get x=2 and x=-2 :)

OpenStudy (welshfella):

probably you haven't come to complex and imaginary numbers yet. They are not real numbers but they do exist in math.

OpenStudy (anonymous):

i have i know what they are...somewhat

OpenStudy (welshfella):

Yes - mathematicians introduced them because some problems could not be solved using real numbers alone.

OpenStudy (welshfella):

Hayleymeyer obviously hasn't been taught them yet.